OpenStudy (anonymous):

calculate the instateneous and the practical buffer capacity in the acid and alkaline direction of 0.01M phosphate buffer, pH 6.8

OpenStudy (aaronq):

it's asking you to find how much acid or base you can add to the buffer specified before the pH changes drastically.

OpenStudy (aaronq):

I would use the Henderson-Hasselbalch equation: \(pH=pKa+log\dfrac{[A^-]}{[HA]}\)

OpenStudy (anonymous):

how can i calculate the practical buffer capacity

OpenStudy (aaronq):

buffer capacity: \(\beta = 2.303 \dfrac{[HA][A^-]}{[HA] + [A^-]}\)

OpenStudy (aaronq):

not sure what they mean by "practical"